OpenStudy (anonymous):

An equation is shown below: 5(2x - 8) + 15 = -15 Write the steps you will use to solve the equation and explain each step.

OpenStudy (anonymous):

I really need help

OpenStudy (anonymous):

ok i got u

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

5(2) is 10x

OpenStudy (anonymous):

5(-8) is -45

OpenStudy (anonymous):

-45+15 is -30

OpenStudy (anonymous):

so -15+30IS????

OpenStudy (anonymous):

I got it from here thanks!

OpenStudy (anonymous):

but here hold

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

$5\left(2x-8\right)+15=-15\quad:\quad x=1$5(2x−8)+15=−15 : x=1 Steps $5\left(2x-8\right)+15=-15$5(2x−8)+15=−15 Show Steps $\mathrm{Expand\:}5\left(2x-8\right)+15:\quad10x-25$Expand 5(2x−8)+15: 10x−25 $10x-25=-15$10x−25=−15 $\mathrm{Add\:}25\mathrm{\:to\:both\:sides}$Add 25 to both sides $10x-25+25=-15+25$10x−25+25=−15+25 $10x=10$10x=10 $\mathrm{Divide\:both\:sides\:by\:}10$Divide both sides by 10 $\frac{10x}{10}=\frac{10}{10}$10x10 =1010 $x=1$x=1

OpenStudy (anonymous):

Is that all or is there more.

OpenStudy (anonymous):

thats all